A car velocity increases from 24 kph to 80 kph while traveling 80 m. Compute its (a) acceleration and (b) the time period during this acceleration.


1
Expert's answer
2021-10-19T15:10:10-0400

(a)

"v^2=v_0^2+2ad,""a=\dfrac{v^2-v_0^2}{2d}=\dfrac{(22.22\ \dfrac{m}{s})^2-(6.67\ \dfrac{m}{s})^2}{2\times80\ m}=2.81\ \dfrac{m}{s^2}."

(b)

"v=v_0+at,""t=\dfrac{v-v_0}{a}=\dfrac{22.22\ \dfrac{m}{s}-6.67\ \dfrac{m}{s}}{2.81\ \dfrac{m}{s^2}}=5.53\ s."